# Neural ODEs for holographic transport models without translation symmetry

### v1

@@ 4.2 Linear axions
The simplest holographic mechanism for breaking translation symmetry is to invoke the linear axion (LA) [59]. Consider the Einstein gravity minimally coupled with two massless scalar fields. Its bulk action is given by
@@
where $X=g^{\mu\nu}X_{\mu\nu}$ and $X_{\mu\nu}=G_{IJ}\partial_{\mu}\chi^{I}\partial_{\nu}\chi^{J}$ with $G_{IJ}=\frac{1}{2}\mathrm{diag}(1,1)$ and $I=1,2$. The Gibbons-Hawking term is same as before but the counterterms are more simple,
$S_{\mathrm{ct}}=\int d^{3}x\sqrt{-\gamma}\left(-4-R+\gamma^{ij}X_{ij}\right).$ (31)
@@
With Eq. (34) and Eq. (38) in hand, we can generate the data. Here we fix the parameter $\beta=1$. Other setting is similar to before. The data $\left[\omega,\eta_{\mathrm{re}}(\omega)\right]$ are plotted in Fig. 4.
Figure 4: The real part of frequency-dependent shear viscosity in the LA model.
The training results are reported in Fig. 5, Fig. 8, and Tab. 1.
Figure 5: The performance of Neural ODEs for learning metric (left) and mass (right) in the LA model.
@@ 4.3 Generalized axions
In the above two models with broken translation symmetry, the mass square of the graviton has only one term. One may wonder if Neural ODEs can learn more general mass function. Here we will study the generalized axion (GA) model [60], which is a productive tool in the field of holographic condensed matters [61], see a recent application in amorphous solids [62]. Of particular interest to us, it can assign more general mass function to the graviton, while remaining analytically tractable.
@@
We will select the function $V(X)$ that has been used in [60]
@@
where $\gamma_{1}$ and $\gamma_{5}$ are two parameters. In terms of the holographic renormalization studied in [63], one can infer that the counterterms are the same as those in the LA model.
@@
Figure 6: The real part of frequency-dependent shear viscosity in the GA model.
Using Eq. (43) and Eq. (47), the data are generated in Fig. 6. Here we have fixed the parameters $\gamma_{1}=\gamma_{5}=1$. We exhibit the training results in Fig. 7, Fig. 8, and Tab. 1.
Figure 7: The performance of Neural ODEs for learning metric (left) and mass (right) in the GA model.
@@ 5 Conclusion and discussion
This work introduces an important physical ingredient in AdS/DL: translation symmetry breaking. Specifically, we extend the data-driven holographic transport model proposed in [34], allowing the machine to learn the black hole metric or the graviton mass from the real part of the shear viscosity. In addition, we use the Neural ODE to learn continuous functions, thereby avoiding the problem of discrete errors. These are the main physical and technical results of this work. Around them, we have the following comments.
@@
Although some assumptions still remain, the current theoretical framework, which mainly consists of the wave equation (3), the regularity at the horizon (7), and the boundary relation (8), is very general. Considering that the calculation of shear viscosity does not depend on the probe limit, this generality is quite rare.
@@
In Tab. 1, it is shown that the Neural ODE can achieve high accuracy for the bulk reconstruction. However, we caution that this is associated with the current model parameters ($\alpha,\beta,\gamma$), which are simply fixed to 1. As these parameters increase, the temperature of black holes decreases. Meanwhile, we observe a rapid degradation in the performance of Neural ODEs, although increasing the data can mitigate this trend [^4] Similar phenomena were previously reported in [35].. In the future, we expect to develop new machine learning techniques or introduce reasonable physical constraints to improve the performance of Neural ODEs at low temperatures.
4) Beyond AdS/CFT
The current theoretical framework necessitates the bulk spacetime to be asymptotic AdS. However, this information is not explicitly fed into the machine. Consequently, it would be plausible to construct more general data-driven holographic models, where the asymptotic geometry could be either Lifshitz or hyperscaling violated. Ultimately, the present work falls within the realm of reverse engineering of AdS/CFT, where the holographic dictionary is given a priori. One can also develop holographic machine learning that is independent of AdS/CFT [67, 64, 65, 66], which explores the emergence of spacetime and gravity from a more fundamental level.
Note added: Following the completion of this work, we became aware of Ref. [68], where the Neural ODE is used to learn the metric from the optical conductivity based on the linear axion model.
@@ Acknowledgments
We thank Xian-Hui Ge, Yan Liu, Yu Tian, and Zhuo-Yu Xian for helpful discussions. SFW was supported by NSFC grants (No.11675097, No.12275166, and No. 12311540141).
@@ Appendix A From shear response to shear viscosity
where $L_{\mathrm{matter}}^{(1)}$ denotes the counterterms contributed by the matter.
We expand the on-shell bulk action, the Gibbons-Hawking term, and the counterterms to the second order of the shear mode,
$\left.S_{\mathrm{bulk}}+S_{\mathrm{GH}}+S_{\mathrm{ct}}\right|_{\mathrm{on-shell}}=\int d^{2}x\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{2}\left[\left(L_{\mathrm{gravity}}^{(2)}+L_{\mathrm{matter}}^{(2)}\right)\bar{h}h-r^{2}\sqrt{\frac{g_{tt}}{g_{rr}}}\bar{h}h^{\prime}\right],$ (A.6)
where $\bar{h}$ has the argument $-\omega$. In the terms $\sim\bar{h}h$, $L_{\mathrm{gravity}}^{(2)}$ can be referred as the gravity contribution and $L_{\mathrm{matter}}^{(2)}$ denotes the contribution from $L_{\mathrm{matter}}^{(1)}$. We find that the former has a universal expression [^5] In deriving Eq. (A.7), the key step is to calculate the second-order variation $\frac{\delta^{2}\int d^{4}x\sqrt{-g}L_{\mathrm{matter}}}{\delta g_{xy}^{2}}=\frac{\sqrt{-g}}{g_{xx}^{2}}\left(\frac{\delta T_{xy}}{\delta g_{xy}}-2\frac{T_{xx}}{g_{xx}}\right).$
$L_{\mathrm{gravity}}^{(2)}=-\frac{r^{2}\omega^{2}}{\sqrt{g_{tt}}}+4r^{2}\sqrt{g_{tt}}-2r\sqrt{\frac{g_{tt}}{g_{rr}}}-\frac{r^{2}g_{tt}^{\prime}}{\sqrt{g_{tt}g_{rr}}},$ (A.7)
while both $L_{\mathrm{gravity}}^{(2)}$ and $L_{\mathrm{matter}}^{(2)}$ generally diverge at the boundary.
Inputting Eq. (A.3) and Eq. (A.4) into Eq. (A.6), we obtain the renormalized action
@@
where $L_{\mathrm{gravity}}^{(3)}$ and $L_{\mathrm{matter}}^{(3)}$ are finite and they are contributed by the gravity and the matter, respectively. One can find that $L_{\mathrm{gravity}}^{(3)}$ is real since it is composed of the real constants in Eq. (A.3). We also argue that $L_{\mathrm{matter}}^{(3)}$ is usually real. This is based on the following facts and assumptions.
1) The counterterms should be the intrinsic scalars defined on the boundary.
2) The induced metric has no $\gamma^{tx}$ and $\gamma^{ty}$ components on the background level.
3) All other fields coupled to $\gamma_{xy}$ are supposed to be time-independent.
4) The gauge symmetry should be respected.
5) The system is supposed to be rotationally invariant.
Using 1)-3), one can see that $L_{\mathrm{matter}}^{(3)}$ cannot have an imaginary part unless the expansion of $L_{\mathrm{matter}}^{(1)}$ contains a term
$L_{\mathrm{im}}\sim\gamma^{tt}\partial_{t}\gamma_{xy}\gamma_{xy}\left(A,\partial A\right)_{t\cdots},$ (A.9)
where $\left(A,\partial A\right)_{t\cdots}$ denotes certain matter field or its spatial derivative with the indices $t\cdots$. The expression of $L_{\mathrm{im}}$ is understood that the subscript $t$ in $\partial_{t}\gamma_{xy}$ must be paired with the subscript $t$ in $\left(A,\partial A\right)_{t\cdots}$ in order to contract with the superscripts of $g^{tt}$. Obviously, $A$ is neither a scalar nor a spinor since they have no spacetime indices. $A$ is not a U(1) vector if 4)-5) are taken into account further. Thus, there is no such term at least in any Einstein-Maxwell-Dirac-scalar models.
To proceed, we extract the retarded correlator from $S_{\mathrm{ren}}$:
@@
Importantly, the last term of Eq. (A.12), which depends on the UV details of specific models, is imaginary.
In view of this, we will consider the Kubo formula for the real part of the frequency-dependent shear viscosity [69, 70, 34]

### v2

@@ 4.2 Linear axions
The simplest holographic mechanism for breaking translation symmetry is to invoke the linear axion (LA) Andrade:1311 . Consider the Einstein gravity minimally coupled with two massless scalar fields. Its bulk action is given by
@@
where $X=g^{\mu\nu}X_{\mu\nu}$ and $X_{\mu\nu}=G_{IJ}\partial_{\mu}\chi^{I}\partial_{\nu}\chi^{J}$ with $G_{IJ}=\frac{1}{2}\mathrm{diag}(1,1)$ and $I=1,2$.
@@
With eq. (32) and eq. (36) in hand, we can generate the data. Here we fix the parameter $\beta=1$. Other setting is similar to before. The data $\left[\omega,\eta_{\mathrm{re}}(\omega)\right]$ are plotted in the left panel of figure 3. The training results are displayed in the middle and right panels of figure 3, as well as in figure 6 and table 1.
Figure 3: Training data and performance of Neural ODEs for the LA model. Left panel: the real part of the frequency-dependent shear viscosity. Middle and right panels: The performance of Neural ODEs in learning the metric and mass.
@@ 4.3 Generalized axions
In the above two models with broken translation symmetry, the mass square of the graviton has only one term. One may wonder if Neural ODEs can learn more general mass function. Here we will study the generalized axion (GA) model Baggioli2014 , which is a productive tool in the field of holographic condensed matters Matteo2021 , see a recent application in amorphous solids Li2108 . Of particular interest to us, it can assign more general mass function to the graviton, while remaining analytically tractable.
@@
We will select the function $V(X)$ that has been used in Baggioli2014
@@
where $\gamma_{1}$ and $\gamma_{5}$ are two parameters.
@@
Using eq. (41) and eq. (45), the data are generated in the left panel of figure 4. Here we have fixed the parameters $\gamma_{1}=\gamma_{5}=1$. We exhibit the training results in middle and right panels of figure 4, as well as in figure 6 and table 1.
Figure 4: Training data and performance of Neural ODEs for the GA model.
@@ 5 Conclusion and discussion
This work introduces an important physical ingredient to AdS/DL: translation symmetry breaking. Specifically, we extend the data-driven holographic transport model proposed in Wu2020 , allowing the machine to learn the black hole metric or the graviton mass from the real part of the shear viscosity [^4] In Wu2020 , the model is subject to the several constraints: minimal coupling of 3+1 dimensional Einstein gravity and matter, homogeneous and isotropic background, vanishing graviton mass, and the relation $\eta=\chi+i\omega r$ on the boundary. In the current work, we allow for a non-zero graviton mass and reduce the boundary relation to its real part.. In addition, we use the Neural ODE to learn continuous functions, thereby avoiding the problem of discretization errors. Around these physical and technical results, we have made the following comments and explorations.
@@
Although some assumptions still remain, the current theoretical framework, which mainly consists of the wave equation (3), the regularity at the horizon (7), and the boundary relation (8), is very general. Considering that the holographic calculation of shear viscosity does not impose the probe limit, this generality is quite rare.
@@
In table 1, it is shown that the Neural ODE can achieve high accuracy for the bulk reconstruction. However, we caution that this is associated with the current model parameters ($\alpha,\beta,\gamma$), which are simply fixed to 1. As these parameters increase, the temperature of black holes decreases. Meanwhile, we observe a rapid degradation in the performance of Neural ODEs [^5] Similar phenomena were previously reported in Li:2022zjc .. To address this issue, we will input the temperature and several inductive biases, the most important of which is to change the loss function. In this way, Neural ODEs can still perform well, see Appendix C.
4) Entanglement entropy
In Hashimoto:2018bnb , after the bulk metric is learned from lattice QCD data, the quark-antiquark potential is calculated using the holographic method. Intriguingly, this potential exhibits both a linear confining part and a Debye screening part. In Appendix D, we conduct a similar exploration.
Using the bulk metric $f(z)$ learned from the shear viscosity $\eta_{\mathrm{re}}(\omega)$, we calculate the derivative $S^{\prime}(l)$, where $S$ is the holographic EE and $l$ is the spatial size of entangling region [^6] We focus on the derivative of EE rather than the EE itself because the former is not sensitive to the UV cutoff Jokela:2020auu ; Jokela2304 ; XWB2023 .. Across a wide range of $l$ spanning several orders of magnitude, we find that the MRE of $S^{\prime}(l)$ is very small. This indicates that there is an intriguing relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$. Note that this relationship may be very complex from the perspective of field theories, partly because $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ belong to two distinct classes of observables: quantum information measures and transport coefficients. Meanwhile, $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ are mediated by the bulk metric, connected via the Neural ODE and the Ryu-Takayanagi (RT) formula Ryu:2006bv , see figure 5. More importantly, the relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ is independent with the probe limit and is not restricted to a specific model. Therefore, we suspect that it may appear in some real-world strongly coupled systems. In particular, given the recent progress in lattice gauge theories with respect to both $S^{\prime}(l)$ and the spectral function related to shear viscosity Jokela2304 ; Altenkort2211 ; Yao2402 , we expect that this conjecture can be verified in the near future.
Figure 5: The schematic diagram for the relationship between $\eta_{\mathrm{re}}(\omega)$ and $S^{\prime}(l)$. At the gravity level, they are connected by the Neural ODE and the RT formula through the bulk metric $f(z)$, which is represented by the straight arrows above. The winding arrow below implies that their relationship may be very complex from the perspective of field theories.
5) Beyond AdS/CFT
The current theoretical framework necessitates that the bulk spacetime be asymptotically AdS. However, we deliberately refrain from explicitly communicating this information to the machine. This intentional omission increases the training difficulty considerably, but the performance of Neural ODEs remains excellent. This suggests the potential for developing a machine-learning framework that can extend beyond asymptotic AdS, capable of accommodating the Lifshitz and hyperscaling-violated geometries, as well as other novel behaviors [^7] By constructing a free bulk theory from the generalized free fields on the boundary, it was found recently that the Sachdev-Ye-Kitaev model retains geometric bulk features beyond its conformal limit Qi2306 . Notably, the curvature of the emergent spacetime diverges near the boundary.. Ultimately, this work falls within the realm of reverse engineering of AdS/CFT, where the holographic dictionary is given a priori. One can also develop holographic machine learning that is independent of AdS/CFT You:2017guh ; Hu:2019nea ; Han:2019wue ; Lam:2021ugb , which explores the emergence of spacetime and gravity from a more fundamental level.
Note added: Following the completion of this work, we became aware of ref. Ahn2401 , where Neural ODEs are used to learn from optical conductivity based on the linear axion model. Encouragingly, it is illustrated that the bulk metric can be reconstructed from the data of the real material UPd2Al3.
@@ Acknowledgments
We thank Xian-Hui Ge, Yan Liu, Zhen-Kang Lu, Yu Tian, and Zhuo-Yu Xian for helpful discussions. SFW was supported by NSFC grants (No.11675097, No.12275166, and No. 12311540141).
@@ Appendix A From shear response to shear viscosity
where $L_{\mathrm{matter}}^{(1)}$ denotes the contribution of the counterterms from the matter. We will expand the on-shell bulk action, the Gibbons-Hawking term, and the counterterms to the second order of the shear mode. Here the key step is to calculate the second-order variation of the matter action $S_{\mathrm{m}}=\int d^{4}x\sqrt{-g}L_{\mathrm{matter}}$ as follows:
$\frac{\delta^{2}S_{\mathrm{m}}}{\delta g_{xy}^{2}}$ $=$ $\frac{1}{2}\left[\frac{\delta S_{\mathrm{m}}}{\delta g^{xx}}\frac{\delta^{2}g^{xx}}{\left(\delta g_{xy}\right)^{2}}+\frac{\delta^{2}S_{\mathrm{m}}}{\left(\delta g^{xx}\right)^{2}}\left(\frac{\delta g^{xx}}{\delta g_{xy}}\right)^{2}+x\leftrightarrow y\right]$ (A.6)
$+\frac{1}{2}\left[\frac{\delta S_{\mathrm{m}}}{\delta g^{xy}}\frac{\delta^{2}g^{xy}}{\left(\delta g_{xy}\right)^{2}}+\frac{\delta^{2}S_{\mathrm{m}}}{\left(\delta g^{xy}\right)^{2}}\left(\frac{\delta g^{xy}}{\delta g_{xy}}\right)^{2}+x\leftrightarrow y\right].$
Since $T_{\mu\nu}=-2\frac{1}{\sqrt{-g}}\frac{\delta S_{\mathrm{m}}}{\delta g^{\mu\nu}}$ and
$g^{xx}=g^{yy}=\frac{g_{xx}}{-g_{xy}^{2}+g_{xx}^{2}},\;g^{xy}=\frac{-g_{xy}}{-g_{xy}^{2}+g_{xx}^{2}},$ (A.7)
one can read
$\frac{\delta^{2}S_{\mathrm{m}}}{\delta g_{xy}^{2}}=-\frac{\sqrt{-g}}{g_{xx}^{2}}\left(\frac{1}{g_{xx}}T_{xx}-\frac{1}{2}\frac{\delta T_{xy}}{\delta g_{xy}}\right).$ (A.8)
Using eq. (A.8) and the Einstein equation for the background metric
$\frac{1}{2}T_{xx}=-3g_{xx}-\frac{g_{xx}g_{rr}^{\prime}g_{tt}^{\prime}}{4g_{rr}^{2}g_{tt}}-\frac{g_{xx}g_{tt}^{\prime 2}}{4g_{rr}g_{tt}^{2}}-\frac{g_{rr}^{\prime}g_{xx}^{\prime}}{4g_{rr}^{2}}+\frac{g_{tt}^{\prime}g_{xx}^{\prime}}{4g_{rr}g_{tt}}-\frac{g_{xx}^{\prime 2}}{4g_{rr}g_{xx}}+\frac{g_{xx}g_{tt}^{\prime\prime}}{2g_{rr}g_{tt}}+\frac{g_{xx}^{\prime\prime}}{2g_{rr}},$ (A.9)
we can perform the expansion and express the result in the frequency space
$\left.S_{\mathrm{bulk}}+S_{\mathrm{GH}}+S_{\mathrm{ct}}\right|_{\mathrm{on-shell}}=\int d^{2}x\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{2}\left[\left(L_{\mathrm{gravity}}^{(2)}+L_{\mathrm{matter}}^{(2)}\right)\bar{h}h-g_{xx}\sqrt{\frac{g_{tt}}{g_{rr}}}\bar{h}h^{\prime}\right],$ (A.10)
where $\bar{h}$ has the argument $-\omega$. In the terms $\sim\bar{h}h$, $L_{\mathrm{gravity}}^{(2)}$ can be referred as the gravity contribution and $L_{\mathrm{matter}}^{(2)}$ denotes the contribution from $L_{\mathrm{matter}}^{(1)}$. One can find that $L_{\mathrm{gravity}}^{(2)}$ has the form
$L_{\mathrm{gravity}}^{(2)}=-\omega^{2}\frac{g_{xx}}{\sqrt{g_{tt}}}+4\sqrt{g_{tt}}g_{xx}-\frac{g_{xx}g_{tt}^{\prime}}{\sqrt{g_{tt}g_{rr}}}-\sqrt{\frac{g_{tt}}{g_{rr}}}g_{xx}^{\prime}.$ (A.11)
Interestingly, it is independent with the energy-momentum tensor and the graviton mass. In contrast, we cannot write down the general form of $L_{\mathrm{matter}}^{(2)}$.
One may notice that $L_{\mathrm{gravity}}^{(2)}$ and $L_{\mathrm{matter}}^{(2)}$ usually diverge at the boundary. To deal with the divergence, we input eq. (A.3) and eq. (A.4) into eq. (A.10), which yields the renormalized action
@@
where $L_{\mathrm{gravity}}^{(3)}$ and $L_{\mathrm{matter}}^{(3)}$ are finite and they are contributed by the gravity and the matter, respectively. In terms of eq. (A.11), it is obvious that $L_{\mathrm{gravity}}^{(3)}$ is real since it is composed of the real constants in eq. (A.3). Moreover, we will assume that $L_{\mathrm{matter}}^{(3)}$ is also real. This assumption is valid for all holographic models in the main text. To prove this, we will list their counterterms explicitly.
1) Massive gravity
The counterterms of massive gravity have been derived in Wu1903 . They can be expressed as
$S_{\mathrm{ct}}=\int d^{3}x\sqrt{-\gamma}\left[-4-R+\frac{1}{2}\alpha e_{1}+\frac{1}{16}\alpha^{2}e_{2}+\frac{\alpha}{4}(2R_{ij}Y^{ij}-e_{1}R)\log r\right],$ (A.13)
from which we calculate
$L_{\mathrm{matter}}^{(2)}=-\frac{1}{8}\alpha\left(\sqrt{g_{tt}g_{xx}}+2\omega^{2}\sqrt{\frac{g_{xx}}{g_{tt}}}\log r\right).$ (A.14)
2) Linear axions
The counterterms of the linear axion model are simple Andrade:1311 :
$S_{\mathrm{ct}}=\int d^{3}x\sqrt{-\gamma}\left(-4-R+\gamma^{ij}X_{ij}\right).$ (A.15)
We can read
$L_{\mathrm{matter}}^{(2)}=-\frac{1}{2}\beta^{2}\sqrt{g_{tt}}.$ (A.16)
3) Generalized axions
In terms of the holographic renormalization studied in Ma2208 , one can infer that the counterterms of the current generalized axion model are given by
$S_{\mathrm{ct}}=\int d^{3}x\sqrt{-\gamma}\left(-4-R+\gamma_{1}\gamma^{ij}X_{ij}\right),$ (A.17)
which indicates
$L_{\mathrm{matter}}^{(2)}=-\frac{1}{2}\gamma_{1}\sqrt{g_{tt}}.$ (A.18)
From (A.14), (A.16), and (A.18), it is obvious that they can only generate real $L_{\mathrm{matter}}^{(3)}$.
Motivated by these explicit examples, let’s move on a more general analysis. Suppose that a counterterm is an intrinsic scalar on the boundary, which consists of the induced metric, matter fields, and their derivatives. To be safe, we will focus on neutral matter fields and impose the parity symmetry [^8] Then the Levi-Civita tensor can be excluded in building the counterterm.. One can see that $L_{\mathrm{matter}}^{(3)}$ cannot have an imaginary part unless the expansion of $L_{\mathrm{matter}}^{(1)}$ contains a term
$L_{\mathrm{im}}\sim\gamma^{tt}\partial_{t}\gamma_{xy}\gamma_{xy}\left(\Psi,\partial\Psi\right)_{t\cdots},$ (A.19)
where $\left(\Psi,\partial\Psi\right)_{t\cdots}$ denotes certain matter field or its spatial derivative with the indices $t\cdots$. The expression of $L_{\mathrm{im}}$ is understood that the subscript $t$ in $\partial_{t}\gamma_{xy}$ must be paired with the subscript $t$ in $\left(\Psi,\partial\Psi\right)_{t\cdots}$ in order to contract with the superscripts of $g^{tt}$. Obviously, $\Psi$ is neither a scalar nor a spinor since they have no spacetime indices. $\Psi$ is also not a U(1) vector: Due to the gauge symmetry, the U(1) vector can only appear in the form of $F_{ij}F^{ij}$, which cannot be paired with $\partial_{t}\gamma_{xy}$. In light of these analysis, one can find that our assumption is valid for all Einstein-Maxwell-Dirac-scalar models with parity symmetry and neutral matters. For more general models, we cannot provide a similar proof by checking eq. (A.19). Nevertheless, keeping in mind the intrinsic structure of counterterms [^9] Even if one can construct a counterterm ansatz that yields $L_{\mathrm{im}}$, this does not necessarily imply that it is a genuine counterterm for a specific model. In Ma2208 , it was found that the exact ansatz of counterterms can be generated from the radial Hamiltonian, indicating the existence of an intrinsic structure of counterterms., it is reasonable to suspect that the current assumption may be valid in a more general situation [^10] For example, consider the famous model of holographic superconductor Hartnoll0803 . The charged complex scalar field $\Phi$ is coupled with the U(1) vector $A$, which in principle might produce: $L_{\mathrm{im}}\sim\gamma^{tt}\partial_{t}\gamma_{xy}\gamma_{xy}\left|(\partial_{t}-iA_{t})\Phi\right|$. However, the matter contribution to the counterterm of this model is just $L_{\mathrm{matter}}^{(1)}\sim\left|\Phi\right|^{2}$.. In fact, we are not aware of any obvious counterexamples.
To proceed, we extract the retarded correlation from $S_{\mathrm{ren}}$:
@@
Importantly, the last term of eq. (A.22), which depends on the UV details of specific models, is purely imaginary.
In view of this, we will consider the Kubo formula for the real part of the frequency-dependent shear viscosity Read1207 ; Wu2015 ; Wu2020